OpenStudy (asylum15):

Solveforx: 3sin(x+25o)=1.8 (00 ≤x≤3600)

OpenStudy (apoorvk):

so, dividing both sides by 3, we get: sin(X+25) = 0.6 find inverse in degrees, then deduct 25 degrees from both sides.

sam (.sam.):

3sin(x+25o)=1.8 divide 3 sin(x+25o)=1.8/3 inverse sin \[x+25=\sin^{-1}(1.8/3)\] find the sin values

OpenStudy (asylum15):

Thanks :)

sam (.sam.):

np

OpenStudy (asylum15):

Are you here Sam?

sam (.sam.):

here? \[x+25=\sin^{-1}(1.8/3)\]

OpenStudy (asylum15):

Yeah, from here? I am not sure of myself, :|

sam (.sam.):

you have a scientific calculator?

OpenStudy (asylum15):

yeah

sam (.sam.):

type in sin^{-1}(1.8)/3 and get the value first

OpenStudy (asylum15):

36.87

sam (.sam.):

ok from the quadrants we only do the sides with positive sign in it